8375 15. 2033 100614 3 Metors Acceptance CEMAC TST Wireless Services CAWE Ford Motor Credit σ) 015177 Coneco Giebal Funding cco) Ford Motor Farge&W wlers Prorty Casualty CTARD 125 Mar 15, 203 005 7.113 Amerada HesCA Soure: The Wall Street Journal, Jouny &,2003 Reprintnd with permissien he annual coupon payment for the Wells Fargo Bond? What is the semi-annual coupon payment for Sprint Capital? What is the maturity date for the Travelers Property Casualty bond? How much will each TPC bond holder receive at the maturity date (assume annual coupon payment)? What is the last traded price for Amerada Hesss bond issue? What is the Yield to Maturity for First Energys bond issue? What is the trading volume in dollars for Albertsons bond issue? How many bonds were transacted? Stocks 3 3988 28.57 Wachova WE 04 27 15 29332 3836 -0.33 41 1750 10.27 Wacknhut Corr WH.1 13 10.65-0.44 6.1 3310 15.30 WadRed A WDR53 25 19 305 20.87-0.2 4 40.70 27.70 Walgreen WAG 5 5 31 58392 3194 049 1 2262 15.80 WallaceCS WCS 66 3O 31 186 21.74 -003 0.1 63.94 4372 WalMart WMT30 6 29 65117 50.46 02 5.4 1515 851 Walterind WLT 12 12 6 746 10.25-003 1 39.98 2780 WashMut WM 1oef 3.0 10 32315 36.30 -0.30 764 516 WashPost & WPO 5.60 7 53 117 754.50 -450 0.4 3015 20.42 WashREIT WRE 41 55 19 794 25.60-0.35 3.6 3998 23.49 WsteConn WC 1 3809 37.22-0.78 9 3110 20.20 WasteMgt WMI 20 17740 2350-017 8 1259 7 WrPkTch PIK What is Walgreens ticker symbol What is the highest that Washington Mutual traded at over the last year? What is the lowest price that Washington Mutual traded at over the last year?
How much did Wachovia pay out in dividends over the past year? What was Wachovias Dividend Yield? According to the stock chart above how much are investors willing to pay for every dollar of Waste Managements earnings per share? What was the volume of shares traded in Walter Industry? What was Washington Posts closing price? What was the closing price the day before?
Answer #1

1. Coupon rate of Wells Fargo = 7.250%

annual Coupon Coupon payment = Par value * Coupon rate = 100* 7.250%/2 = 7.25

2.  

Semi annual Coupon rate of Wells Fargo = 8.375%

annual Coupon Coupon payment = Par value * Coupon rate = 100* 7.375%/2 = 4.1875 or 4.19

3. Maturity date of TPC bond = March 15,2013
Amount received at end of maturity date = Coupon + Par value = 100* 5% + 100 = 105

4. Amerada Hess's Bond issue = 100.145